OpenStudy (anonymous):

The Ka for HCN is 6.2 x 10^-10 Calculate: a) [H3O+] b) the pH and the % dissociation for a 0.50 mol/L HCN solution

OpenStudy (mayankdevnani):

@peekaboopork i think the concentration of HCN is given in question ! because for (a) part :- \[\huge \bf [H_3O^{+}]=\sqrt{K_a \times C}\] where, \[\large \bf c=concentration~of~molecule\]

OpenStudy (anonymous):

Yes. From there, you can find pH by using \(-log[H_3O^+]\). The \(\%~dissociation=\large \frac{[H^+]}{[HCN]}\)